OpenStudy (anonymous):

Find sin θ if cot θ = - 4 and cos θ < 0.

OpenStudy (anonymous):

Use this identity \[\csc ^{2} \theta - \cot ^{2} \theta = 1\] solve for csc theta and use \[\csc \theta = 1/ sin \theta\]